OpenStudy (anonymous):

Please help! Line AB contains points A (−1, 3) and B (5, 3). Line AB has a slope that is zero undefined positive negative

OpenStudy (texaschic101):

to find the slope, use the slope formula : slope = (y2 - y1) / (x2 - x1) (-1,3) x1 = -1 and y1 = 3 (5,3) x2 = 5 and y2 = 3 now we sub slope = (3 - 3) / (5 - 1) slope = 0/4 ok...remember this...when the 0 is in the numerator, it is a zero slope.It is also a horizontal line. If the 0 would have been in the denominator, it would have been an undefined slope and would have been a vertical line.

OpenStudy (anonymous):

Thanks a millions! This helped me so much you don't even know!

OpenStudy (texaschic101):

glad to help :)